OpenStudy (anonymous):

A student attempts to solve for b in the equation below. Which statement best applies to the work shown below?

OpenStudy (anonymous):

1/2b-2=1 1/2b*2/1-2=1 b=2=1 b-2+2=1+2 b=3 A. The student incorrectly applied the multiplication property of equality before the addition property of equality. B. The student did not apply the addition property of equality to isolate b. C. The student did not apply the multiplication property of equality to isolate b. D. The student correctly solved the two-step equation.

OpenStudy (anonymous):

I know its not d but i cant figure out which other it is

OpenStudy (jdoe0001):

\(\bf \cfrac{1}{2}b-2=1\\ \quad \\ \cfrac{1}{2}b\times \color{red}{\cfrac{2}{1}}-2=1\\\quad \\ b-2=1\\\quad \\ b-2+2=1+2\\\quad \\ b=3\)

OpenStudy (jdoe0001):

see anything wrong there?

OpenStudy (anonymous):

what will the answer be im so lost